# Concentration of discrepancy–based approximate Bayesian computation via Rademacher complexity

### v4

@@ 3.1 Limiting behavior with fixed $\varepsilon_{n}=\varepsilon$ and $n\rightarrow\infty$
Current abc theory for the setting $\varepsilon_{n}=\varepsilon$ and $n\rightarrow\infty$, studies convergence of the discrepancy–based abc posterior
$\smash{\pi^{(\varepsilon)}_{n}(\theta)}\propto\pi(\theta)\int_{\mathcal{Y}^{n}}\mathds{1}\{\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\hat{\mu}_{y_{1:n}})\leq\varepsilon\}\mu_{\theta}^{n}(dz_{1:n}),$
@@
under the key assumption that $\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\hat{\mu}_{y_{1:n}})\rightarrow\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})$ almost surely as $n\rightarrow\infty$ (see, e.g., Jiang et al., 2018, Theorem 1). When such a condition is met, it can be shown that $\lim_{n\rightarrow\infty}\smash{\pi^{(\varepsilon)}_{n}(\theta)}\propto\pi(\theta)\mathds{1}\{\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})\leq\varepsilon\}$. This means that the limiting abc–posterior coincides with the prior constrained within the support $\{\theta:\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})\leq\varepsilon\}$, where $\varepsilon$ is the same threshold employed for $\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\hat{\mu}_{y_{1:n}})$.
@@
Although this is a key result, verifying the convergence assumption behind the above finding can be difficult for several discrepancies (Jiang et al., 2018). More crucially, such a theory fails to provide informative results in those situations where $\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\hat{\mu}_{y_{1:n}})$ is not guaranteed to converge to $\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})$. In fact, under such a setting, the asymptotic properties of discrepancy–based abc posteriors have been overlooked to date, and it is not even clear whether a well–defined $\smash{\pi^{(\varepsilon)}_{n}}$ exists in the large–data limit. Indeed, for specific choices of the discrepancy and threshold, the abc posterior may not be well–defined even for fixed $n$. For example, when $\mathcal{D}_{\mathfrak{F}}$ is the tv distance (see Example A.1.1), then $\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\hat{\mu}_{y_{1:n}})=1$, almost surely, whenever $z_{1:n}$ and $y_{1:n}$ are from two continuous distributions $\mu_{\theta}$ and $\mu^{*}$, respectively. Hence, for any $\varepsilon<1$, $\smash{\pi^{(\varepsilon)}_{n}}$ is not defined, even if $\mu_{\theta}$ and $\mu^{*}$ are not mutually singular. Note that, as clarified in Example A.2.1, the Rademacher complexity of the tv distance is always bounded away from zero in such a setting. In this case and, more generally, for ips discrepancies whose $\mathfrak{R}_{n}(\mathfrak{F})$ does not vanish to zero, Lemma 2.1 together with the triangle inequality in (2.1) still allow to derive new informative upper and lower bounds for limiting acceptance probabilities which further guarantee that $\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})$ can be bounded from above, for a sufficiently large $n$, by $\varepsilon+4\mathfrak{R}_{n}(\mathfrak{F})$, even without assuming that $\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\hat{\mu}_{y_{1:n}})\rightarrow\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})$ almost surely for $n\rightarrow\infty$ as in the current convergence theory. Theorem 3.1 formalizes this intuition for the whole class of abc posteriors arising from discrepancies in the ips family.
@@
Let $\mathcal{D}_{\mathfrak{F}}$ be an ips as in Definition 2.1. Moreover, assume (I), (III), and let $c_{\mathfrak{F}}=4\lim\sup\mathfrak{R}_{n}(\mathfrak{F})$. Then, the acceptance probability $p_{n}=\int_{\Theta}\int_{\mathcal{Y}^{n}}\mathds{1}\{\mathcal{D}(\hat{\mu}_{z_{1:n}},\hat{\mu}_{y_{1:n}})\leq\varepsilon\}\mu_{\theta}^{n}(dz_{1:n})\pi(d\theta)$ of the rejection–based abc routine with discrepancy $\mathcal{D}_{\mathfrak{F}}$ satisfies
$\qquad\pi\{\theta:\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})\leq\varepsilon-c_{\mathfrak{F}}\}+o(1)\leq p_{n}\leq\pi\{\theta:\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})\leq\varepsilon+c_{\mathfrak{F}}\}+o(1),$ (3.1)
almost surely with respect to $y_{1:n}\stackrel{{\scriptstyle\mbox{\scriptsize i.i.d.}}}{{\sim}}\mu^{*}$, as $n\rightarrow\infty$, for any fixed $\varepsilon>0$.
In particular, whenever $\varepsilon>\tilde{\varepsilon}+c_{\mathfrak{F}}$, with $\tilde{\varepsilon}=\inf\{\varepsilon>0:\smash{\pi\{\theta:\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})\leq\tilde{\varepsilon}\}}>0\}$, the abc acceptance probability $p_{n}$ is bounded away from 00, for $n$ large enough. This implies that the abc posterior $\pi_{n}^{(\varepsilon)}$ is always well–defined for a large–enough $n$ and its support is almost surely included asymptotically within the set $\{\theta:\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})\leq\varepsilon+4\mathfrak{R}_{n}(\mathfrak{F})\}$. Namely
@@
Theorem 3.1 clarifies that, even when relaxing the convergence assumptions behind current theory, the newly–introduced Rademacher complexity framework still allows to obtain guarantees on the acceptance probabilities, existence and limiting support of the abc posterior. Such a relaxation crucially highlights the key role of the richness of $\mathfrak{F}$ in driving the asymptotic behavior of $\pi_{n}^{(\varepsilon)}$. As previously discussed, the higher is $\mathfrak{R}_{n}(\mathfrak{F})$ the lower are the guarantees that $\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\hat{\mu}_{y_{1:n}})$ is close enough to $\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})$ in the large data limit. In fact, recalling Section 2, when $\mathfrak{R}_{n}(\mathfrak{F})$ does not vanish with $n$ there are no guarantees that the convergence of $\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\hat{\mu}_{y_{1:n}})$ to $\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})$ assumed in the current literature is verified. Nonetheless, the inequalities in Lemma 2.1 still yield informative results which correctly translate these arguments into an inflation of the superset that contains the support of the limiting abc posterior, and in a more conservative choice for the threshold $\varepsilon$ to ensure that $\pi_{n}^{(\varepsilon)}$ is well–defined for large $n$. Notice that in Theorem 3.1 the dependence on the dimensions $d$ and $p$ is embedded within $\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})$; see Equation (6) in the Appendix of Chérief-Abdellatif and Alquier (2022) for an example making such a dependence more explicit.
Recall that, for $\pi_{n}^{(\varepsilon)}$ to be well–defined in the limit (with a fixed threshold $\varepsilon>0$), the acceptance probability $p_{n}$ needs to be bounded away from zero. According to Theorem 3.1, as $n\rightarrow\infty$ this is guaranteed whenever $\varepsilon>\tilde{\varepsilon}+c_{\mathfrak{F}}$. For instance, recalling Section 2, in misspecified models $\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})\geq\varepsilon^{*}$. As a consequence, $\varepsilon>c_{\mathfrak{F}}$ is not sufficient to guarantee that $\pi\{\theta:\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})\leq\varepsilon-c_{\mathfrak{F}}\}$ in (3.1) is bounded away from zero. In fact, when $c_{\mathfrak{F}}<\varepsilon<c_{\mathfrak{F}}+\varepsilon^{*}$ the lower bound in (3.1) is exactly 00. For this event to happen, $\varepsilon-c_{\mathfrak{F}}$ needs to define a ball with radius not lower than the one to which the prior is already guaranteed to assign strictly positive mass. This implies $\varepsilon>\tilde{\varepsilon}+c_{\mathfrak{F}}$, with $c_{\mathfrak{F}}=4\lim\sup\mathfrak{R}_{n}(\mathfrak{F})$.
The results in Theorem 3.1 also clarify that, when $\mathfrak{R}_{n}(\mathfrak{F})\rightarrow 0$ as $n\rightarrow\infty$, stronger and uniform convergence statements can be derived, both in correctly–specified and in misspecified models. More specifically, in this setting, the acceptance probability $p_{n}$ in Theorem 3.1 approaches $\pi\{\theta:\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})\leq\varepsilon\}$ in the limit. Moreover, the abc posterior is well–defined for any $\varepsilon>\tilde{\varepsilon}$ and its asymptotic support has the same $\varepsilon$–control as the one established by the abc procedure on the discrepancy among the empirical distributions. These results allow to establish the almost sure convergence of the abc posterior stated in Corollary 3.1.
Corollary 3.1.
Under the conditions in Theorem 3.1, if also $\mathfrak{R}_{n}(\mathfrak{F})\rightarrow 0$ as $n\rightarrow\infty$ (i.e., Assumption (IV) is satisfied), then, for any $\varepsilon>\tilde{\varepsilon}$, it holds that
$\pi_{n}^{(\varepsilon)}(\theta)\rightarrow\pi(\theta\mid\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})\leq\varepsilon)\propto\pi(\theta)\mathds{1}\left\{\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})\leq\varepsilon\right\},$ (3.2)
@@
Corollary 3.1 clarifies that to obtain within the whole ips class a convergence result in line with those stated in the current theory it is not necessary to assume that $\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\hat{\mu}_{y_{1:n}})$ converges almost surely to $\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})$. In fact, it is sufficient to verify that $\mathfrak{R}_{n}(\mathfrak{F})\rightarrow 0$ as $n\rightarrow\infty$. Crucially, this condition is imposed directly on the selected discrepancy within the ips class, rather than on the model or the underlying data generating process and, therefore, can be constructively verified for all the discrepancies in Examples 2.12.3. For instance, while Jiang et al. (2018) is unable to provide conclusive guidelines on whether (3.2) holds under Wasserstein distance and mmd, our Corollary 3.1 can be directly applied to prove convergence for both these divergences, leveraging the results and discussions in Section 3.3.
@@
Unlike current theory, Corollary 3.1 accounts also for misspecified models. In fact, the limiting pseudo–posterior in (3.2) is well–defined only when $\varepsilon>~{}\tilde{\varepsilon}$. For example, when $\mu^{*}$ is not within $\{\mu_{\theta}:\theta\in\Theta\subseteq\mathbb{R}^{p}\}$, then $\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})\geq\varepsilon^{*}$. Hence, for any $\varepsilon\leq\varepsilon^{*}$, we have that $\mathds{1}\{\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})\leq\varepsilon\}=0$ for any $\theta\in\Theta$ and the limiting pseudo–posterior is not well–defined.
@@
Motivated by the convergence result in Corollary 3.1, the theory in Section 3.2 provides an in–depth study of the concentration properties for the abc posterior when also the tolerance threshold progressively shrinks. As is clear from (3.2), by employing a vanishing threshold might guarantee that, in the limit, the abc posterior $\pi_{n}^{(\varepsilon)}$ concentrates all its mass around $\mu^{*}$, in correctly specified settings, or around the distribution $\mu_{\theta^{*}}$ closest to $\mu^{*}$, among those in $\{\mu_{\theta}:\theta\in\Theta\subseteq\mathbb{R}^{p}\}$ for misspecified regimes. These results and the associated rates of concentration are provided in Section 3.2 by leveraging, again, Rademacher complexity theory.

### v5

@@ 3.1 Limiting behavior with fixed $\varepsilon_{n}=\varepsilon$ and $n\rightarrow\infty$
Current abc theory for the setting $\varepsilon_{n}=\varepsilon$ fixed and $n\rightarrow\infty$, studies convergence of the discrepancy-based abc posterior
${\pi^{(\varepsilon)}_{n}(\theta)}\propto\pi(\theta)\int_{\mathcal{Y}^{n}}\mathds{1}\{\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\hat{\mu}_{y_{1:n}})\leq\varepsilon\}\mu_{\theta}^{n}(dz_{1:n}),$
@@
under the key assumption that $\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\hat{\mu}_{y_{1:n}})\rightarrow\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})$ almost surely as $n\rightarrow\infty$ (see, e.g., Jiang, Wu and Wong, 2018, Theorem 1). When such a condition is met, it can be shown that $\lim_{n\rightarrow\infty}{\pi^{(\varepsilon)}_{n}(\theta)}\propto\pi(\theta)\mathds{1}\{\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})\leq\varepsilon\}$. This means that the limiting abc posterior coincides with the prior constrained within the support $\{\theta:\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})\leq\varepsilon\}$, where $\varepsilon$ is the same threshold employed for $\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\hat{\mu}_{y_{1:n}})$.
@@
Although this is a key result, verifying the convergence assumption behind the above finding can be difficult for several discrepancies (e.g., Jiang, Wu and Wong, 2018). More crucially, such theory fails to provide informative results in those situations where $\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\hat{\mu}_{y_{1:n}})$ is not guaranteed to converge to $\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})$. In fact, under this setting, the asymptotic properties of discrepancy-based abc posteriors have been overlooked, and it is not even clear whether a well-defined ${\pi^{(\varepsilon)}_{n}}$ exists in the limit. Indeed, for specific choices of the discrepancy and threshold, the abc posterior may not be well-defined even for fixed $n$. For example, when $\mathcal{D}_{\mathfrak{F}}$ is the tv distance (see Appendix A in the Supplementary Material), then $\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\hat{\mu}_{y_{1:n}})=1$, almost surely, whenever $z_{1:n}$ and $y_{1:n}$ are from two continuous distributions $\mu_{\theta}$ and $\mu^{*}$, respectively. Hence, for any $\varepsilon<1$, the abc posterior is not defined, even if $\mu_{\theta}$ and $\mu^{*}$ are not mutually singular. Note that, as clarified in Appendix A, the Rademacher complexity of the tv distance is always bounded away from zero in such a setting. In this case and, more generally, for ips discrepancies whose $\mathfrak{R}_{n}(\mathfrak{F})$ does not vanish to zero, Lemma 2.6 together with the triangle inequality in (1) still allow to derive new informative upper and lower bounds for limiting acceptance probabilities which further guarantee that $\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})$ can be bounded from above, for a sufficiently large $n$, by $\varepsilon+4\mathfrak{R}_{n}(\mathfrak{F})$, even without assuming that $\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\hat{\mu}_{y_{1:n}})\rightarrow\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})$ almost surely for $n\rightarrow\infty$, as in the current convergence theory. Theorem 3.1 formalizes this intuition for the whole class of abc posteriors arising from discrepancies in the ips family.
@@
Let $\mathcal{D}_{\mathfrak{F}}$ be an ips as in Definition 2.1. Moreover, assume (I), (III), and let $c_{\mathfrak{F}}=4\lim\sup\mathfrak{R}_{n}(\mathfrak{F})$. Then, the acceptance probability $p_{n}=\int_{\Theta}\int_{\mathcal{Y}^{n}}\mathds{1}\{\mathcal{D}(\hat{\mu}_{z_{1:n}},\hat{\mu}_{y_{1:n}})\leq\varepsilon\}\mu_{\theta}^{n}(dz_{1:n})\pi(d\theta)$ of rejection abc with discrepancy $\mathcal{D}_{\mathfrak{F}}$ satisfies, for any fixed $\varepsilon>0$,
$\qquad\pi\{\theta:\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})\leq\varepsilon-c_{\mathfrak{F}}\}-e_{n}\leq p_{n}\leq\pi\{\theta:\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})\leq\varepsilon+c_{\mathfrak{F}}\}+e_{n},$ (3)
almost surely with respect to $y_{1:n}\stackrel{{\scriptstyle\mbox{\scriptsize i.i.d.}}}{{\sim}}\mu^{*}$, as $n\rightarrow\infty$, where $e_{n}=\exp(-\sqrt{n}/(2b^{2}))=o(1)$.
In particular, whenever $\varepsilon>\tilde{\varepsilon}+c_{\mathfrak{F}}$, with $\tilde{\varepsilon}=\inf\{\epsilon>0:{\pi\{\theta:\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})\leq\epsilon\}}>0\}$, the probability $p_{n}$ is bounded away from 00, for $n$ large enough. This implies that, for every fixed $\varepsilon>\tilde{\varepsilon}+c_{\mathfrak{F}}$, the abc posterior is always well-defined for a large-enough $n$ and its support is almost surely included asymptotically in the set $\{\theta:\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})\leq\varepsilon+4\mathfrak{R}_{n}(\mathfrak{F})\}$. Namely,
@@
for any fixed $\varepsilon>\tilde{\varepsilon}+c_{\mathfrak{F}}$, almost surely with respect to $y_{1:n}\stackrel{{\scriptstyle\mbox{\scriptsize i.i.d.}}}{{\sim}}\mu^{*}$, as $n\rightarrow\infty$.
Theorem 3.1 clarifies that, even when relaxing the assumptions behind current theory, the Rademacher complexity framework still allows to obtain guarantees on the acceptance probabilities, existence and limiting support of the abc posterior. Such a relaxation highlights the key role of the richness of $\mathfrak{F}$ in driving these properties. As previously discussed, the higher is $\mathfrak{R}_{n}(\mathfrak{F})$ the lower are the guarantees that $\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\hat{\mu}_{y_{1:n}})$ is close enough to $\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})$ in the large data limit. In fact, recalling Section 2, when $\mathfrak{R}_{n}(\mathfrak{F})$ does not vanish with $n$ there are no guarantees that the convergence of $\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\hat{\mu}_{y_{1:n}})$ to $\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})$ assumed in the literature is verified. Nonetheless, the inequalities in Lemma 2.6 still yield informative results that translate these arguments into an inflation of the superset containing the support of the limiting abc posterior, and in a more conservative choice for the threshold $\varepsilon$ to ensure that such a posterior is well-defined for large $n$. Notice that in Theorem 3.1 the dependence on the dimension of the parameter and data space is embedded within $\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})$; see Equation (6) in the Appendix of Chérief-Abdellatif and Alquier (2022) for an example making this dependence explicit.
Recall that, for the abc posterior to be well-defined in the limit (with a fixed threshold $\varepsilon$), the acceptance probability $p_{n}$ needs to be bounded away from zero. According to Theorem 3.1, as $n\rightarrow\infty$ this is guaranteed whenever $\varepsilon>\tilde{\varepsilon}+c_{\mathfrak{F}}$. For instance, recalling Section 2, in misspecified models $\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})\geq\varepsilon^{*}$. As a consequence, $\varepsilon>c_{\mathfrak{F}}$ is not sufficient to guarantee that $\pi\{\theta:\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})\leq\varepsilon-c_{\mathfrak{F}}\}$ in (3) is bounded away from zero (in fact, when $c_{\mathfrak{F}}<\varepsilon<c_{\mathfrak{F}}+\varepsilon^{*}$ the lower bound in (3) approaches zero from below). To achieve this result, $\varepsilon-c_{\mathfrak{F}}$ needs to define a ball with radius not lower than the one to which the prior is already guaranteed to assign strictly positive mass. This implies $\varepsilon>\tilde{\varepsilon}+c_{\mathfrak{F}}$, with $c_{\mathfrak{F}}=4\lim\sup\mathfrak{R}_{n}(\mathfrak{F})$.
The results in Theorem 3.1 further clarify that, when $\mathfrak{R}_{n}(\mathfrak{F})\rightarrow 0$ as $n\rightarrow\infty$, stronger and uniform convergence statements can be derived, both in correctly-specified and in misspecified models. More specifically, in this setting, the acceptance probability $p_{n}$ in Theorem 3.1 approaches $\pi\{\theta:\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})\leq\varepsilon\}$ in the limit. Moreover, the abc posterior is well-defined for any $\varepsilon>\tilde{\varepsilon}$ and its asymptotic support has the same $\varepsilon$-control as the one established by the abc procedure on the discrepancy among the empirical distributions. These results allow to establish the almost sure convergence of the abc posterior stated in Corollary 3.2.
Corollary 3.2.
Under the conditions in Theorem 3.1, if also $\mathfrak{R}_{n}(\mathfrak{F})\rightarrow 0$ as $n\rightarrow\infty$ (i.e., Assumption (IV) is satisfied), then, for any fixed $\varepsilon>\tilde{\varepsilon}$, it holds that
$\pi_{n}^{(\varepsilon)}(\theta)\rightarrow\pi(\theta\mid\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})\leq\varepsilon)\propto\pi(\theta)\mathds{1}\left\{\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})\leq\varepsilon\right\},$ (4)
@@
Corollary 3.2 clarifies that to obtain within the whole ips class a convergence result in line with those stated in the current theory it is not necessary to assume that $\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\hat{\mu}_{y_{1:n}})$ converges almost surely to $\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})$. In fact, it is sufficient to verify that $\mathfrak{R}_{n}(\mathfrak{F})\rightarrow 0$ as $n\rightarrow\infty$. Crucially, this condition is imposed directly on the selected discrepancy within the ips class, rather than on the model or the underlying data generating process and, therefore, can be constructively verified for all the discrepancies in Examples 2.22.4. For instance, while Jiang, Wu and Wong (2018) is unable to provide conclusive guidelines on whether (4) holds under Wasserstein distance and mmd, our Corollary 3.2 can be directly applied to prove convergence for both these divergences, leveraging the results and discussions in Section 3.3.
@@
Corollary 3.2 accounts also for misspecified models. In fact, the limiting pseudo-posterior in (4) is well-defined only when $\varepsilon>\tilde{\varepsilon}$. For example, when $\mu^{*}$ is not within $\{\mu_{\theta}:\theta\in\Theta\subseteq\mathbb{R}^{p}\}$, then $\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})\geq\varepsilon^{*}$. Hence, for any $\varepsilon\leq\varepsilon^{*}$, we have that $\mathds{1}\{\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})\leq\varepsilon\}=0$ for any $\theta\in\Theta$ and the limiting pseudo-posterior is not well-defined.
@@
Motivated by the convergence result in Corollary 3.2, the theory in Section 3.2 provides an in-depth study of the concentration properties for the abc posterior when the tolerance threshold progressively shrinks. As is clear from (4), employing a vanishing threshold might guarantee that, in the limit, the abc posterior concentrates all its mass around $\mu^{*}$, in correctly specified settings, or around the distribution $\mu_{\theta^{*}}$ closest to $\mu^{*}$, among those in $\{\mu_{\theta}:\theta\in\Theta\subseteq\mathbb{R}^{p}\}$ for misspecified regimes. These results and the associated rates of concentration are provided in Section 3.2 by leveraging, again, Rademacher complexity theory.